A tie on a vertical pressing machine is 150cm long and 5cm square in cross section. Determine the stress and extension produced in the bar when it is subjected to an axial load of 150kN. Take E = 2.1*10^5N/mm^2


1
Expert's answer
2021-02-01T09:28:17-0500

stress=force/area

f=150,000N

area =15,000*50,000m

150000/0.00075=

stress =50,000N/mm2


extension(strain)=stress/elasticity

E = 2.1*10^5N/mm^2

stress =50,000n/m2

= 50,000N/mm2 / 2.1*10^5N/mm2

= 2.38*109 N/mm2